OpenStudy (anonymous):

Solve for c. (1/4)c + 6 = -1

OpenStudy (anonymous):

Help please

imqwerty (imqwerty):

1 Simplify brackets (1/4)c+6=−1 2 Simplify (1/4)c to c/4 c/4+6=−1 3 Multiply both sides by 4 c+24=−4 4 Subtract 24 from both sides c=−4−24 5 Simplify −4−24 to −28 c=−28